# Multi-Label Bayesian Active Learning with Inter-Label Relationships

### v1

@@ Abstract
The primary challenge of multi-label active learning, differing it from multi-class active learning, lies in assessing the informativeness of an indefinite number of labels while also accounting for the inherited label correlation. Existing studies either require substantial computational resources to leverage correlations or fail to fully explore label dependencies. Additionally, real-world scenarios often require addressing intrinsic biases stemming from imbalanced data distributions. In this paper we proposes an new multi-label active learning strategy to address both challenges. Our method incorporates progressively updated positive and negative correlation matrices to capture co-occurrence and disjoint relationships within the label space of annotated samples, enabling a holistic assessment of uncertainty rather than treating labels as isolated elements. Furthermore, alongside diversity, our model employs ensemble pseudo labeling and beta scoring rules to address data imbalances. Extensive experiments on four realistic datasets demonstrate that our strategy consistently achieves more reliable and superior performance, compared to several established methods.
Multi-Label Bayesian Active Learning with Inter-Label Relationships
Yuanyuan Qi, Jueqing Lu, Xiaohao Yang, Joanne Enticott, Lan Duthanks: Corresponding author. Monash University, {yuanyuan.qi, jueqing.lu, xiaohao.yang, joanne.enticott, lan.du}@monash.edu
@@ 1 Introduction
In recent years, extensive machine learning models and algorithms have been developed to deal with the exponential growth of real-world data. However, the significant mismatch between the rapid increase in data and the slow pace of manual data annotation underscores the imperative of active learning (AL) Liu et al. (2021); Xie et al. (2022). Multi-label active learning (MLAL), which considers the co-occurrence of labels and is more aligned with real-world applications, has been explored in different domains, including text classification Han et al. (2024); Kang et al. (2020), medical imaging Huang et al. (2024); Simao et al. (2023), remote sensing Möllenbrok et al. (2023); Möllenbrok and Demir (2023), and so on. The multi-label task, due to the complexity of label-wise correlation and imbalanced data distribution, remains a vital task to comprehensively examine the data, thus necessitating the development of effective query strategies Siméoni et al. (2020); Dor et al. (2020).
@@
To deal with the multi-label issue in active learning, earlier approaches usually transform it into multiple binary classification tasks, known as binary relevance (BR) which sums the informativeness evaluated for each individual labels to obtain the final acquisition score Zheng et al. (2021); Wang et al. (2022). However, these approaches overlook the potential correlation of labels, such as their co-occurrence, which should be factored into the overall information assessment Zhang et al. (2021); Min et al. (2022). Consequently, the information inherent in the label correlation of the queried samples may not be fully explored.
@@
Some recent works have employed co-occurrence and label correlation matrices to model these inherent label relationships Su et al. (2023). However, while the positive correlations, indicating strong co-occurrence between labels, have been included, few studies have explored negative correlations where labels are mutually exclusive and do not appear together. Moreover, asymmetric label-wise correlations, where one label frequently appears with another label without a reciprocal relationship, remains under-explored. This also includes the hierarchical structure of the label set, where node labels inherently belong to and serve as subsets of their corresponding root labels. The selection of overlapping labels, due to the hierarchical nature, affects the diversity of the strategy, consequently, its overall outcome Nakano et al. (2020). Furthermore, due to the high imbalance ratios in real-world datasets, addressing data imbalance to maintain consistent performance across different datasets highlights the critical importance of MLAL tasks Chen et al. (2022); Arens et al. (2024).
@@
Considering label co-occurrence and data imbalance, we propose a new MLAL framework, named multi-label CoRrelation-Aware active learning with Beta scoring rules (CRAB) in this paper. By incorporating the Beta scoring rules to deal with data imbalance and the expected loss reduction framework to select the most informative data instance, we introduce dynamic positive and negative correlation matrices to handle the distinct and asymmetric label correlation within a Bayesian framework. This approach demonstrates robust and outstanding performance on four benchmark datasets for multi-label active learning.
@@ 2 Related Work
Active learning involves selecting the most informative data from the unlabeled pool for annotation, thereby reducing the required training data while maintaining comparable performance. Two mainstream AL query strategies are uncertainty-based and diversity-based approaches Ren et al. (2021); the former concentrates on informative measurement at the sample-level Abdar et al. (2021), while the latter emphasizes data distribution Kim et al. (2023). To quantify the sample uncertainty, methods such as proper scoring rules Tan et al. (2024), Dirichlet distribution Hemmer et al. (2022), and Gaussian Process Shi et al. (2021) can be used to estimate the sample informativeness. By aligning the prior and posterior distributions with model output observations, these models can effectively capture the uncertainty of each prediction.
@@
However, focusing exclusively on uncertainty can introduce bias in sampling (i.e., selecting near-identical instances, thus wasting the annotation budget), which may lead to sub-optimal performance Prabhu et al. (2019). Incorporating diversity into the sampling process offers an alternative approach to enhancing generalization Buchert et al. (2023). Huang and Zhou Huang and Zhou (2013) utilized label cardinality inconsistency to exploit uncertainty and integrated it with the diversity-based sampling during data acquisition. PLVI-CE leverages average posterior probability discrepancy to measure data diversity and prediction inconsistency to assess uncertainty, thus enhancing model generalization with limited annotated instances Gu et al. (2023). Recently, Tan et al. Tan et al. (2024) proposed BESRA which uses the strictly proper scoring rules. Its acquisition function combines beta-scoring rules and k-means clustering to enhance diversity, while the Beta scoring rules also address data imbalance common in multi-label datasets. Built on top of BESRA, our framework further takes into account label correlation in the acquisition function.
@@
Research in active learning has gradually paid attention to the label correlation in MLAL. In light of the specific characteristics of graph data, DAMAL incorporates class-label interactions using a graph-based ranking approach, where edge weights are defined as the cosine similarity between latent features, thus quantifying the graph’s informativeness in relation to label correlation Mahapatra et al. (2024); Lu et al. (2020). To address the uncertainty in feature correlations within standard data, Shi et al. Shi et al. (2021) integrated a Gaussian process with a Bernoulli Mixture model to model correlation through the covariance matrix. Correlation matrix-based weighted uncertainty, typically derived through co-occurrence or label similarity analysis, is commonly used to query the most informative label pairs by capturing the inter-label influence during label selection Gong and Zhai (2021); Su et al. (2023). Han et al. Han et al. (2024) propose a two-stage sample acquisition strategy, called ALMuLa-mix, utilizing inconsistency to capture label correlations with novel features as the first stage and employing the class frequency at the second stage to ensure inter-class diversity.
@@
3 Correlation-aware Active Learning
@@
3.1 Preliminary: Beta Scoring Rules
Monte Carlo estimation of error reduction Roy and McCallum (2001) involves using the Monte Carlo approach to estimate the expected reduction in error resulting from the labeling of a query. Inspired by this, BEMPS Tan et al. (2023) improves these formulations by employing the proper scoring rules Gneiting and Raftery (2007). Rather than estimate the error, the scoring rule provides a summary measure of predictive probability, which computes the positive oriented rewards (i.e., utilities) that a classier maximizes Gneiting and Raftery (2007).
@@ 3.1 Preliminary: Beta Scoring Rules
$\quad\quad[S(P(x,\theta),y)-S(P(x),y)]\big{]}$ (1)
@@
Eqs (1) and (2) show the core concept of expected increase in score when querying Tan et al. (2023). $S$ denotes the scoring function that evaluates the predictive probability distribution on an event, i.e., predicting $y^{\prime}$, given $x^{\prime}$. $Q(L)$ represents the mean proper scoring rule of the predictive probabilities obtained using Bayesian estimation based on the current labeled dataset $L$. The term $\Delta Q(x|L)$ denotes the increment in the score resulting from acquiring the label of a sample $x$, drawn from the unlabeled data pool $U$. And $x$ denotes a selected unlabeled anchor point from $x^{\prime}$ for assessment. The value of $\Delta Q(x|L)$ is then used to select the sample leading to a large increment in the score or reward. Since the label of unlabeled points $x$ and $x^{\prime}$ is unknown, we derive the $P(y^{\prime}|L+\{x,y\},x^{\prime})$ by calculating the posterior distribution of the ensemble models using Eq. (3).
@@
To address the issue of imbalanced label distribution in text datasets, BESRA Tan et al. (2024) leverage Beta family Buja et al. (2005), which generalizes the logarithmic score and the brier score, or the other desired cost-weighted scoring rule. The equation below illustrates the proper scoring rules $\mathcal{L}$ of a predictive distribution $p$ given the expected value $y^{k}$, where $y^{k}$ represents the label for class $k$. $\mathcal{L}(1|p)$ and $\mathcal{L}(0|p)$ represent the partial losses when $p$ is been classified as 1 and 0, respectively.
$S^{k}_{BR}(p,y^{k})\quad$ $=\quad\mathcal{L}(y^{k}|p)$
$=\quad y^{k}\mathcal{L}(1|p)+(1-y^{k})\mathcal{L}(0|p)$ (5)
@@
By leveraging $I_{x}(\alpha,\beta)$, the Incomplete Beta Function, the closed form of the Beta distribution is obtained for $\alpha,\beta>0$. When $\alpha=\beta=0$, the scoring becomes log-loss, and when $\alpha=\beta=1$, the scoring rule will transform to squared error losses. By adjusting the value of $\alpha$ and $\beta$, our model can effectively handle scenarios with diverse data distributions. In our research, we employ the greedy search result of BESRA as the parameter for scoring, where the $\alpha=0.1,\beta=3$ Tan et al. (2024).
@@ 3.2 Correlation Matrix Construction
In multi-label scenario, a single sample often has more than one label, and label with relatively small semantic distances frequently appear simultaneously. To leverage inherent relationships within the label space to assist in acquisition process, our framework maintains two dynamic correlation matrices: a co-occurrence matrix representing positive correlations, and a non-co-occurrence matrix representing negative correlations between labels. Both matrices are updated after each acquisition iteration with newly annotated instances.
By discovering the pattern of the occurrence between labels, we aim to quantify the informativeness considering influence of correlation, and make the query process diversity while take into account the imbalanced data distribution.
3.2.1 Positive correlation matrix:
The positive correlation matrix is derived from the label-wise dependence matrix $A(m,n)$, which represent the dependence of one label’s presence on another. This can be implemented via Eq. (8). Specifically, $\sum_{i=1}^{L}N(y_{i}^{m}=+1,y_{i}^{n}=+1)$ refers to the count of labeled instances where labels $m$ and $n$ occur simultaneously in the label set. $P(y^{m}|y^{n})$ gives the likelihood of label $m$ occurring when label $n$ is present. When $m=n$, $A(m,n)$ equals 1. The value of $A(m,n)$ reflects the probability of one label’s existence to another.
@@ 3.2.1 Positive correlation matrix:
For example, if label $m$ is the subset of the label $n$, the presence of label $m$ definitely implies the presence of label $n$, while the converse does not hold. In this way, $A(m,n)$ will equal 1, and $A(n,m)$ generally smaller than $A(m,n)$. Thereby, the positive correlation matrix is able to describe the pattern of the co-occurrence between labels and reflects the asymmetric correlation between labels, including hierarchical relationships.
3.2.2 Negative correlation matrix
Despite the positive correlations, negative correlations between labels have rarely been addressed in previous research. However, in real-world scenarios, negative correlations are instrumental in enabling the model differentiate between mutually exclusive classes and contribute to more accurate decisions by clarifying the model’s decision boundaries Yang et al. (2024).
To effectively model these negative correlations, we construct an updated non-co-occurrence matrix, $NegA(m,n)$ where $m\neq n$, as defined in Eq. (9). This matrix quantifies the confidence in the presence of label $m$ given that label $n$ is present. Specifically, $\sum_{i=1}^{L}N(y_{i}^{m}=-1,y_{i}^{n}=+1)$ represents the count where labels $m$ and $n$ do not co-occur. If label $m$ and $n$ occur simultaneously, then both $NegA(m,n)$ and $NegA(n,m)$ are set to 1. While in most cases, $NegA(m,n)$ dose not equal to $NegA(n,m)$, as it depends on the occurrence frequency of the baseline label. This asymmetry allows the matrix to capture the nuanced conditional negative relationships, and provide a new perspective towards the label dependencies.
@@ 3.3 Correlation-based Sampling
Refine the unlabeled pool to ensure that selected instances concentrate on specific representative criteria is a common strategy in active learning Kang et al. (2020). However, current research predominantly based on informativeness analysis, neglecting the critical role of data correlation in MLAL.
To address this limitation and provide more representative and evenly distribution samples for continuous process, our model refines the unlabeled pool from three perspectives based on the correlation properties. And the pseudo label $y^{*}$, obtained by averaging the prediction result of ensemble models through Eq. (10), is used for the following correlation-based sampling.
$y_{i}^{*}=\mathbb{I}[P(y_{i}\mid x_{i},L)>0.5]$ (10)
$P(y_{i}\mid x_{i},L)$ $=\textstyle\int_{\theta}P(y_{i}|x_{i},\theta)p(\theta\mid L)$
$\approx\textstyle\sum^{E}_{e=1}P(y_{i}|x_{i},\theta_{e})/E$ (11)

### v2

@@ Abstract
The primary challenge of multi-label active learning, differing it from multi-class active learning, lies in assessing the informativeness of an indefinite number of labels while also accounting for the inherited label correlation. Existing studies either require substantial computational resources to leverage correlations or fail to fully explore label dependencies. Additionally, real-world scenarios often require addressing intrinsic biases stemming from imbalanced data distributions. In this paper, we propose a new multi-label active learning strategy to address both challenges. Our method incorporates progressively updated positive and negative correlation matrices to capture co-occurrence and disjoint relationships within the label space of annotated samples, enabling a holistic assessment of uncertainty rather than treating labels as isolated elements. Furthermore, alongside diversity, our model employs ensemble pseudo labeling and beta scoring rules to address data imbalances. Extensive experiments on four realistic datasets demonstrate that our strategy consistently achieves more reliable and superior performance, compared to several established methods.
@@ 1 Introduction
In recent years, extensive machine learning models and algorithms have been developed to deal with the exponential growth of real-world data. However, the significant mismatch between the rapid increase in data and the slow pace of manual data annotation underscores the imperative of active learning (AL) [Liu et al., 2021, Xie et al., 2022]. Multi-label active learning (MLAL), which considers the co-occurrence of labels and is more aligned with real-world applications, has been explored in different domains, including text classification [Han et al., 2024, Kang et al., 2020], medical imaging [Huang et al., 2024, Simao et al., 2023], remote sensing [Möllenbrok et al., 2023, Möllenbrok and Demir, 2023], and so on. The multi-label task, due to the complexity of label-wise correlation and imbalanced data distribution, remains a vital task to comprehensively examine the data, thus necessitating the development of effective query strategies [Siméoni et al., 2020, Dor et al., 2020].
@@
To deal with the multi-label issue in active learning, earlier approaches usually transform it into multiple binary classification tasks, known as binary relevance (BR) which sums the informativeness evaluated for each individual label to obtain the final acquisition score [Zheng et al., 2021, Wang et al., 2022]. However, these approaches overlook the potential correlation of labels, such as their co-occurrence, which should be factored into the overall information assessment [Zhang et al., 2021, Min et al., 2022]. Consequently, the information inherent in the label correlation of the queried samples may not be fully explored.
@@
Some recent works have employed co-occurrence and label correlation matrices to model these inherent label relationships [Su et al., 2023]. However, while the positive correlations, indicating strong co-occurrence between labels, have been included, few studies have explored negative correlations where labels are mutually exclusive and do not appear together. Moreover, asymmetric label-wise correlations, where one label frequently appears with another label without a reciprocal relationship, remains under-explored. This also includes the hierarchical structure of the label set, where node labels inherently belong to and serve as subsets of their corresponding root labels. The selection of overlapping labels, due to the hierarchical nature, affects the diversity of the strategy, consequently, its overall outcome [Nakano et al., 2020]. Furthermore, due to the high imbalance ratios in real-world datasets, addressing data imbalance to maintain consistent performance across different datasets highlights the critical importance of MLAL tasks [Chen et al., 2022, Arens et al., 2024].
@@
Considering label co-occurrence and data imbalance, we propose a new MLAL framework, named multi-label CoRrelation-Aware active learning with Beta scoring rules (CRAB)  [^1] Our code is publicly available at https://github.com/qijindou/CRAB. in this paper. By incorporating the Beta scoring rules to deal with data imbalance and the expected loss reduction framework to select the most informative data instance, we introduce dynamic positive and negative correlation matrices to handle the distinct and asymmetric label correlation within a Bayesian framework. This approach demonstrates robust and outstanding performance on four benchmark datasets for multi-label active learning.
@@ 2 Related Work
Active learning involves selecting the most informative data from the unlabeled pool for annotation, thereby reducing the required training data while maintaining comparable performance. Two mainstream AL query strategies are uncertainty-based and diversity-based approaches [Ren et al., 2021]; the former concentrates on informative measurement at the sample-level [Abdar et al., 2021], while the latter emphasizes data distribution [Kim et al., 2023]. To quantify the sample uncertainty, methods such as proper scoring rules [Tan et al., 2024], Dirichlet distribution [Hemmer et al., 2022], and Gaussian Process [Shi et al., 2021] can be used to estimate the sample informativeness. By aligning the prior and posterior distributions with model output observations, these models can effectively capture the uncertainty.
@@
However, focusing exclusively on uncertainty can introduce bias in sampling (i.e., selecting near-identical instances, thus wasting the annotation budget), which may lead to sub-optimal performance [Prabhu et al., 2019]. Incorporating diversity into the sampling process offers an alternative approach to enhancing generalization [Buchert et al., 2023]. Huang and Zhou [2013] utilized label cardinality inconsistency to exploit uncertainty and integrated it with the diversity-based sampling. PLVI-CE leverages average posterior probability discrepancy to measure data diversity and prediction inconsistency to assess uncertainty, thus enhancing model generalization with limited annotated instances [Gu et al., 2023]. Recently, Tan et al. [2024] proposed BESRA which uses the strictly proper scoring rules. Its acquisition function combines beta-scoring rules and k-means clustering to enhance diversity, while the Beta scoring rules also address data imbalance common in multi-label datasets. Inspired by BESRA, our framework further takes into account label correlation in the acquisition function.
@@
Research in active learning has gradually paid attention to the label correlation in MLAL. In light of the specific characteristics of graph data, DAMAL incorporates class-label interactions using a graph-based ranking approach, where edge weights are defined as the cosine similarity between latent features, thus quantifying the graph’s informativeness in relation to label correlation [Mahapatra et al., 2024, Lu et al., 2020]. To address the uncertainty in feature correlations within standard data, Shi et al. [2021] integrated a Gaussian process with a Bernoulli Mixture model to model correlation through the covariance matrix. Correlation matrix-based weighted uncertainty, typically derived through co-occurrence or label similarity analysis, is commonly used to query the most informative label pairs by capturing the inter-label influence during label selection [Gong and Zhai, 2021, Su et al., 2023]. Han et al. [2024] propose a two-stage sample acquisition strategy, called ALMuLa-mix, utilizing inconsistency to capture label correlations with novel features as the first stage and employing the class frequency at the second stage to ensure inter-class diversity.
@@
3 Correlation-aware Multi-label Active Learning
@@
3.1 Preliminaries
To address the multi-label active learning problem, most studies decompose it into multiple binary classification tasks, aggregating individual label scores instead of assessing the entire instance holistically. This approach is formulated in Eq. (1), where $S_{BR}$ represents a scoring function that measures the informativeness (e.g., uncertainty score) of individual samples, and $S_{BR}^{k}$ denotes the score with respect to each label.
$S_{BR}(p,y)=\sum^{K}_{k=1}S^{k}_{BR}(p,y^{k})$ $=\sum^{K}_{k=1}\mathcal{L}(y^{k}|p)$ (1)
Monte Carlo estimation offers a probabilistic framework for computing acquisition scores by incorporating randomness to account for variability. Monte Carlo-based error reduction estimation [Roy and McCallum, 2001] utilizes the Monte Carlo approach to approximate the expected reduction in error resulting from the labeling of a given sample. However, while Monte Carlo-based methods estimate expected improvement in model performance, they do not assess the quality of probabilistic outputs. Rather than estimating error, proper scoring rules provide a summary measure of predictive probability, computing the positive-oriented rewards (i.e., utilities) that a classifier seeks to maximize [Gneiting and Raftery, 2007]. Eq. (2) and (3) show the core concept of expected increase in score when querying.
@@ 3.1 Preliminaries
$\quad\quad[S(P(\cdot|x,\theta),y)-S(P(\cdot|x),y)]\big{]}$ (2)
@@
$S$ denotes the scoring function that evaluates predictive probability distribution on an event, i.e., predicting $y^{\prime}$, given $x^{\prime}$. $Q(L)$ represents the mean proper scoring rule of the predictive probabilities obtained using Bayesian estimation based on the current labeled dataset $L$. The term $\Delta Q(x|L)$ denotes the increment in the score resulting from acquiring the label of a sample $x$, drawn from the unlabeled data pool $U$. And $x$ is the point to be acquired, $x^{\prime}$ denotes the selected unlabeled anchor point for assessment. The value of $\Delta Q(x|L)$ is then used to select sample leading to a large increment in the score or reward. Since the label of unlabeled points $x$ and $x^{\prime}$ is unknown, we derive $P(y^{\prime}|L,(x,y),x^{\prime})$ by calculating the posterior distribution of the ensemble models using Eq. (4).
@@
Beta family [Buja et al., 2005], which generalizes the logarithmic score and the brier score, or the other desired cost-weighted scoring rule, is able to address the issue of imbalanced label distribution in multi-label learning. The equation below illustrates the proper scoring rules $\mathcal{L}$ of a predictive distribution $p$ given the expected value $y^{k}$, where $y^{k}$ represents the label for class $k$. $\mathcal{L}(1|p)$ and $\mathcal{L}(0|p)$ represent the partial losses when $p$ is been classified as 1 and 0, respectively.
$S^{k}_{BR}(p,y^{k})\quad=\quad y^{k}\mathcal{L}(1|p)+(1-y^{k})\mathcal{L}(0|p)$ (6)
@@
By leveraging $I_{x}(\alpha,\beta)$, the Incomplete Beta Function, the closed form of the Beta distribution is obtained for $\alpha,\beta>0$. When $\alpha=\beta=0$, the scoring becomes log-loss, and when $\alpha=\beta=1$, the scoring rule will transform to squared error losses. By adjusting the value of $\alpha$ and $\beta$, our model can effectively handle scenarios with diverse data distributions. In our research, we employ the greedy search result of BESRA as the parameter for scoring, where the $\alpha=0.1,\beta=3$ [Tan et al., 2024].
@@ 3.2 Correlation Matrix Construction
In multi-label scenario, a single sample often has more than one label, and label with relatively small semantic distances frequently appear simultaneously. To leverage inherent relationships within the label space to assist in acquisition process, our framework maintains two dynamic matrices: a co-occurrence matrix representing positive correlations, and a anti-correlation matrix representing negative correlations between labels. Both matrices are updated after each acquisition iteration with newly annotated instances. By discovering the pattern of the occurrence between labels, we aim to quantify the informativeness considering influence of correlation, and maintain the diversity while take into account the imbalanced data distribution.
Positive correlation matrix: The positive correlation matrix $A$ is constructed based on the label-wise dependence. $A$ is a $K\times K$ two dimensional matrix, where each element $A(m,n)$ quantifies the dependency of the presence of label $m$ on label $n$. This dependency is formally computed using Eq. (9). Specifically, $\sum_{i=1}^{L}N(y_{i}^{m}=+1,y_{i}^{n}=+1)$ refers to the count of labeled instances in which both labels $m$ and $n$ appear simultaneously. $P(y^{m}|y^{n})$ gives the likelihood of label $m$ occurring when label $n$ is present. When $m=n$, $A(m,n)$ equals 1. The value of $A(m,n)$ reflects the probability of one label’s existence conditioned on another.
@@
The positive correlation matrix is constructed to characterize the pattern of the co-occurrence between labels and capture the asymmetric correlation between labels, including hierarchical relationships.
Negative correlation matrix: Despite the positive correlations, negative correlations between labels have rarely been addressed in previous research. However, in real-world scenarios, negative correlations are instrumental in enabling the model differentiate between mutually exclusive classes and contribute to more accurate decisions by clarifying the model’s decision boundaries [Yang et al., 2024].
To effectively model these negative correlations, we construct an updated anti-correlation matrix, $NegA$. Maintaining the same format as the positive correlation matrix $A$, $NegA$ is a $K\times K$ two dimensional matrix, where each element $Neg(m,n)$ quantifies the confidence in the absence of label $m$ given the presence of label $n$, as defined in Eq. (10). Specifically, $\sum_{i=1}^{L}N(y_{i}^{m}=-1,y_{i}^{n}=+1)$ represents the count of instances where labels $m$ and $n$ do not co-occur. This asymmetry allows the matrix to capture the nuanced conditional negative relationships, and provide a new perspective towards the label dependencies.
@@ 3.3 Correlation-based Sampling
Refining the unlabeled pool to ensure that selected instances concentrate on specific representative criteria is a common strategy in active learning [Kang et al., 2020]. However, current research predominantly based on informativeness analysis, neglecting the critical role of data correlation in MLAL. To address this limitation and provide more representative and evenly distribution samples for continuous process, our model refines the unlabeled pool from three perspectives based on the correlation properties to generate a subset to be used in acquisition. And the pseudo label $y^{*}$, obtained by averaging the prediction result of ensemble models through Eq. (11), is used for the following correlation-based sampling.
$y^{*}$ $=\mathbb{I}[P(y\mid x,L)>0.5]$ (11)
$P(y\mid x,L)$ $=\textstyle\int_{\theta}P(y|x,\theta)p(\theta\mid L)$
$\approx\textstyle\sum^{E}_{e=1}P(y|x,\theta_{e})/E$ (12)